OpenStudy (anonymous):

Will fan + medal Question Below: What is the slope of the line that passes through the point (6,4) and (3,8)? A. 3/4 B. -3/4 C. 4/3 D. -4/3 Thank you for helping! :)

OpenStudy (anonymous):

Picture Reference: https://gyazo.com/ad970580d4605696c15b15e7d0bd547e

OpenStudy (trojanpoem):

dy/dx = y_{2} - y{1} / x_{2} - x_{1} ?

OpenStudy (anonymous):

I would say B... -3/4.. but my math might be wrong

OpenStudy (anonymous):

just do rise over run

OpenStudy (anonymous):

So like, 4 - 8... 6 - 3?

OpenStudy (anonymous):

-4/3 maybe

OpenStudy (anonymous):

-3/4 maybe acurate.

OpenStudy (anonymous):

6 - 3 = 3 ))) ((( 4 -8 = -4... So the answer would be -4/3? I would say 3/-4 but there is no option for that.

OpenStudy (trojanpoem):

8 - 4 / 3-6 = 4/-3 = -4/3

OpenStudy (trojanpoem):

Why maybe :) ? Just plug numbers then you're sure.

OpenStudy (anonymous):

Oh Ok I see thank you! @TrojanPoem @ObeyFCS Sorry I sound really stupid lately. Studying up for a big test on Friday, not the best at Linear Equations and Inequalities...

OpenStudy (trojanpoem):

@DatChinookGuy , You aren't stupid. If you had been, you would have just passed by the question. Good luck at your test, A+